# Time-Dependent Black Hole Lensing and the Unified Weak-to-Strong Deflection Framework

### v1

@@ II.2 Observer tetrad and gauge
$e^{\mu}{}_{\hat{\phi}}$ $=\frac{1}{r_{o}\sin\theta_{o}}\,\delta^{\mu}{}_{\phi}$, (9)
with $f$ given in Eq. (1) (here $\theta_{o}$ is the polar angle of the observer; the $\hat{\phi}-$leg contains $\sin\theta_{o}$ as usual). These vectors satisfy
$g_{\mu\nu}(x_{o})\,e^{\mu}{}_{\hat{a}}\,e^{\nu}{}_{\hat{b}}=\eta_{\hat{a}\hat{b}}+\mathcal{O}(\varepsilon^{2}),$ (10)
where $x_{o}^{\mu}$ denotes the observer event and $\eta_{\hat{a}\hat{b}}=\mathrm{diag}(-1,1,1,1)$. The $\mathcal{O}(\varepsilon)$ corrections implicit in Eq. (10) are uniquely fixed by enforcing orthonormality with the full metric of Eq. (3); their explicit forms are deferred to Appendix A.
@@
$X=-\frac{r_{o}\,p^{\hat{\phi}}}{p^{\hat{0}}},\qquad Y=\frac{r_{o}\,p^{\hat{\theta}}}{p^{\hat{0}}},\qquad b\equiv\sqrt{X^{2}+Y^{2}},$ (13)
which defines the impact parameter $b$ measured on the screen. Eq. (13) is equivalent to the common celestial-sphere parametrization by small angles $(\theta_{X},\theta_{Y})=(X/r_{o},Y/r_{o})$ at large $r_{o}$. We adopt the convention that +X+X points toward decreasing $\phi$ (hence the minus sign in $X$), and +Y+Y toward increasing $\theta$.
Figure 1: Geometry and observer-screen schematic. The Schwarzschild background with horizon at $r=2M$ (filled disk) and photon sphere at $r_{\rm ph}=3M$ (dashed circle) is shown together with a near-critical null trajectory with closest approach $r_{0}(b)$. A static observer at $r_{o}\gg M$ carries the tetrad of Eq. (9) and measures screen coordinates $(X,Y)$ via Eq. (13). The critical impact parameter $b_{c}=3\sqrt{3}\,M$ is indicated. This setup underlies the background bending integral in Eq. (17) and the time-dependent Born map (Eqs. (19)-(24)).
@@
Because the perturbation is axisymmetric, we may, without loss of generality, take $\theta_{o}=\pi/2$ for most calculations. At zeroth order in $\varepsilon$ the conserved energy $E=-p_{t}$ and angular momentum $L=p_{\phi}$ of the photon in the Schwarzschild background Eq. (1) satisfy
$\frac{X}{r_{o}}\simeq-\frac{L}{E\,r_{o}\sin\theta_{o}},\quad\frac{Y}{r_{o}}\simeq\frac{p_{\theta}}{E\,r_{o}},\quad b=\frac{L}{E}+\mathcal{O}\,\left(\frac{M}{r_{o}}\right),$ (14)
so that $b$ reduces to the usual ratio $L/E$ in the asymptotically flat region. In later sections, we shall describe deflection through a two-dimensional vector $\bm{\alpha}(b,t_{o})$ defined on the screen; its weak- and strong-field forms follow from the master relation introduced in Eq. (1) of Sec. III.
@@
$h_{\mu\nu}\rightarrow h_{\mu\nu}-\nabla_{\mu}\xi_{\nu}-\nabla_{\nu}\xi_{\mu},$ (15)
and on tensorial quantities derived from $h_{\mu\nu}$ accordingly. We implement two safeguards to ensure that screen observables built from Eq. (13) are gauge-independent at $\mathcal{O}(\varepsilon)$. First, all dynamical content of the even-parity perturbation is encoded in the Zerilli gauge-invariant master field $\Psi_{Z}$ of Eqs. (6)-(8), and the reconstruction used in Appendix A expresses $p^{\hat{a}}$ corrections in terms of $\Psi_{Z}$ and its derivatives only. Second, the tetrad is fixed operationally by the normalization and orthogonality conditions in Eq. (10) at the physical observer event $x_{o}^{\mu}$. With these choices, any change induced by Eq. (15) in the local frame is a pure $\mathcal{O}(\varepsilon)$ Lorentz transformation that leaves $(X,Y)$ invariant at that order. Concretely, one finds that the gauge variation of the screen map is a boundary term proportional to $\xi^{\mu}$ evaluated at $r_{o}$, and our asymptotic conditions impose $\xi^{\mu}=\mathcal{O}(1/r_{o})$, hence
$\delta_{\xi}X=\delta_{\xi}Y=0+\mathcal{O}\,\left(\frac{\varepsilon}{r_{o}}\right).$ (16)
Eq. (16) guarantees that the centroid motion, multi-image separations, and photon-ring modulations defined later are bona fide observables at first order.
Finally, we denote by $t_{o}$ the Schwarzschild time coordinate of the detection event $x_{o}^{\mu}$, and by $\lambda_{o}$ the corresponding affine parameter value on the photon worldline. The association $t\mapsto t_{o}$ defines the phase with which the QNM signal in Eqs. (6)-(8) is sampled by the observer and will be the sole time variable entering the deflection $\bm{\alpha}(b,t_{o})$ in Secs. III-V.
@@ III Unified Born Integral and Weak-Field Expansion
We now derive a single, time-dependent line-of-sight expression for the lensing deflection that is valid to first order in the perturbation amplitude $\varepsilon$ of Eq. (3) while keeping the Schwarzschild background of Eq. (1) exact. The resulting “master Born integral” will be evaluated along the unperturbed null geodesic with impact parameter $b$ and will furnish, by asymptotic expansion, both the far-field Born law (Sec. III.2) and the time-dependent strong-deflection form (Sec. IV). All observable angles are defined on the screen introduced in Sec. II.2 and are therefore gauge-safe to $\mathcal{O}(\varepsilon)$.
@@ III.1 Master time-dependent Born integral
Let $x_{0}^{\mu}(\lambda;b)$ be the null geodesic of the Schwarzschild background (Eq. (1)) with conserved energy $E=-p_{0t}$ and angular momentum $L=p_{0\phi}$, and let $p_{0}^{\mu}(\lambda)=dx_{0}^{\mu}/d\lambda$ be its four-momentum. The impact parameter measured at infinity is $b=L/E$ to leading order, consistent with Eq. (14). The background bending angle $\alpha_{0}(b)$ is given exactly by the standard integral Weinberg (1972)
$\alpha_{0}(b)=2\int_{r_{0}(b)}^{\infty}\frac{dr}{r^{2}}\,\left[\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}\right]^{-1/2}-\pi,$ (17)
@@
We switch on the first-order, axisymmetric, even-parity QNM perturbation of Eq. (3), encoded by the Zerilli master field $\Psi_{Z}$ that satisfies Eqs. (6)-(8). In the “Born” or single-scattering approximation on the curved background, the photon follows $x_{0}^{\mu}(\lambda;b)$ while its momentum acquires an $\mathcal{O}(\varepsilon)$ correction from the perturbed Hamiltonian $H=(1/2)g^{\mu\nu}p_{\mu}p_{\nu}=0$. Linearizing Hamilton’s equation $dp_{\mu}/d\lambda=-\partial_{\mu}H$ about the background solution and evaluating the source on $x_{0}^{\mu}$ and $p_{0}^{\mu}$ yields
$\delta p_{\mu}(\lambda_{o};b)=-\frac{1}{2}\int_{\lambda_{s}}^{\lambda_{o}}\,d\lambda\,\partial_{\mu}h_{\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}(\lambda)\,p_{0}^{\beta}(\lambda),$ (18)
where $\lambda_{o}$ and $\lambda_{s}$ denote, respectively, the values at the observer and the source, and indices are raised/lowered with the background metric. Eq. (18) is the covariant Born response for the photon momentum on the fixed Schwarzschild trajectory.
Projecting $\delta p_{\mu}$ on the observer tetrad of Eq. (9) and converting to screen angles via Eq. (13) gives the master deflection map
$\bm{\alpha}(b,t_{o})=\bm{\alpha}_{0}(b)+\varepsilon\,\delta\bm{\alpha}(b,t_{o}),$ (19)
with $\bm{\alpha}_{0}(b)$ determined by Eq. (17), and
$\delta\alpha_{A}(b,t_{o})=\int_{\lambda_{s}}^{\lambda_{o}}\,d\lambda\,\mathcal{K}_{A}\,\left[x_{0}(\lambda);p_{0}(\lambda)\right],\quad A\in{X,Y}.$ (20)
@@
$\times\partial_{\mu}h_{\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}(\lambda)\,p_{0}^{\beta}(\lambda),$ (21)
where $e^{\nu}{}_{\hat{0}}$ is the time leg of the observer tetrad Eq. (9), $\Pi_{A}{}^{\ \mu}$ projects onto the $(X,Y)$ screen directions at the observer Eq. (13), and hats denote tetrad components at $x_{o}^{\mu}$. The above projector form is one convenient representative; Appendix B shows its equivalence to a manifestly covariant $\delta\Gamma$ line integral with cancelling boundary terms under our asymptotic conditions. Different but equivalent kernel representatives follow from integrating Eq. (18) by parts and using the background geodesic equation; all yield the same $\delta\bm{\alpha}$ at $\mathcal{O}(\varepsilon)$. Note that the screen deflection arises from varying $X$ and $Y$ in Eq. (13), yielding
$\delta\alpha_{A}=\frac{1}{p_{\hat{0}}^{0}}\,\Pi_{A}{}^{\hat{b}}\,\delta p_{\hat{b}}$ (22)
The QNM time dependence in Eqs. (6)-(8) implies that, along the background ray, the perturbation takes the form
@@
$\omega$ $=\omega_{R}+i\omega_{I}\quad\text{for}\quad\omega_{I}<0,$ (23)
where $t_{0}(\lambda)$ is the Schwarzschild time along $x_{0}^{\mu}(\lambda)$. Writing $t_{0}(\lambda)=t_{o}-\tau(\lambda)$ with $\tau(\lambda)\geq 0$ the light-travel delay between the encounter point and the observer event, Eq. (20) can be cast as
@@
$\mathcal{A}_{A}(b,\omega)=\int_{\lambda_{s}}^{\lambda_{o}}\,d\lambda\,\tilde{\mathcal{K}}_{A}\,\left[r(\lambda),\theta(\lambda)\right]\,e^{i\omega\,\tau(\lambda)},$ (24)
where $\tilde{\mathcal{K}}_{A}$ is the complex kernel obtained from $\mathcal{K}_{A}$ after factoring out the harmonic time dependence (the explicit form follows from Appendix B together with the RWZ reconstruction in Appendix A). Eq. (24) shows that the measured deflection inherits the single QNM frequency $\omega$; its amplitude $\mathcal{A}_{A}$ depends on $b$ only through the unperturbed trajectory and the delay profile $\tau(\lambda)$.
Eqs. (19)-(24) constitute the master, time-dependent Born formalism. Two regimes will be obtained from the same expression. In Sec. III.2 we expand Eq. (24) for $M/b\ll 1$ to recover the far-field Born law and the centroid wobble. In Sec. IV we analyze the contribution from segments of the trajectory that probe $r\approx 3M$, showing that Eq. (24) reproduces the logarithmic strong-deflection behavior once $b\to b_{c}^{+}=3\sqrt{3}M$, with the time dependence inherited from the same phase factor.
@@ III.2 Far-field (Born) expansion and centroid wobble
We now evaluate Eqs. (19)-(24) in the weak-deflection domain $M/b\ll 1$, keeping the Schwarzschild background exact along the ray but expanding the integrals in powers of $M/b$. The background bending reduces to Einstein (2005); Weinberg (1972); Will (2014)
$\alpha_{0}(b)=\frac{4M}{b}+\mathcal{O}\,\left(\frac{M^{2}}{b^{2}}\right),$ (25)
as obtained from Eq. (17). For the perturbation, we use Eq. (24) with the unperturbed trajectory approximated by a straight line of impact parameter $b$ through the asymptotically flat region, so that
$r(\lambda)\simeq\sqrt{b^{2}+z^{2}},\qquad t_{0}(\lambda)=t_{o}-\tau(\lambda),\qquad\tau(\lambda)\simeq z,$ (26)
where $z$ is the affine proxy ($z=0$ at the observer) for the longitudinal distance along the ray and $t_{o}$ is the detection time
In the far zone the even-parity $\ell=2$ QNM behaves as an outgoing quadrupolar wave with $h_{\mu\nu}\sim r^{-1}e^{-i\omega(t-r_{*})}$, hence $\partial_{\mu}h_{\alpha\beta}\sim r^{-2}\,e^{-i\omega(t-r_{*})}$. Substituting this falloff into Eq. (20) or, equivalently, Eq. (24), we obtain the leading Born correction
$\delta\alpha_{A}(b,t_{o})=\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{c_{A}(\omega)}{b}\,\mathcal{F}_{A}(\omega b)\right\}+\mathcal{O}\,\left(\frac{M}{b^{2}}\right),\qquad A\in{X,Y},$ (27)
@@
$\mathcal{F}_{A}(\omega b)\equiv\int_{-\infty}^{+\infty}\,\frac{dz}{b}\,\widehat{\mathcal{K}}_{A}\,\left(\frac{\sqrt{b^{2}+z^{2}}}{b}\right)\,e^{i\omega z}$ (28)
is a dimensionless Fresnel-type factor encoding the finite-frequency sampling of the QNM along the nearly straight path. The explicit form of $\widehat{\mathcal{K}}_{A}$ is given in Appendix B. Two limits of Eq. (27) are especially transparent:
$\delta\alpha_{A}(b,t_{o})\sim\begin{cases}\displaystyle\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{\tilde{c}_{A}(\omega)}{b}\right\},&|\omega|,b\ll 1,\\ \displaystyle\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{\hat{c}_{A}(\omega)}{b}\,e^{-|\mathrm{Im}\omega|\,b}\right\},&|\omega|,b\gg 1,\end{cases}$ (29)
with $\tilde{c}_{A},\hat{c}_{A}$ calculable constants. Thus the leading time-dependent correction decays as $1/b$ and inherits the QNM phase $e^{-i\omega t_{o}}$, in agreement with the structure of Eq. (24).
Collecting Eqs. (25) and (27), the far-field deflection vector on the screen reads

### v2

@@ II.2 Observer tetrad and gauge
$e^{\mu}{}_{\hat{\phi}}$ $=\frac{1}{r_{o}\sin\theta_{o}}\,\delta^{\mu}{}_{\phi}$, (II.9)
with $f$ given in Eq. (II.1) (here $\theta_{o}$ is the polar angle of the observer; the $\hat{\phi}-$leg contains $\sin\theta_{o}$ as usual). These vectors satisfy [15, 35]
$g_{\mu\nu}(x_{o})\,e^{\mu}{}_{\hat{a}}\,e^{\nu}{}_{\hat{b}}=\eta_{\hat{a}\hat{b}}+\mathcal{O}(\varepsilon^{2}),$ (II.10)
where $x_{o}^{\mu}$ denotes the observer event and $\eta_{\hat{a}\hat{b}}=\mathrm{diag}(-1,1,1,1)$. The $\mathcal{O}(\varepsilon)$ corrections implicit in Eq. (II.10) are uniquely fixed by enforcing orthonormality with the full metric of Eq. (II.3); their explicit forms are deferred to Appendix A.
@@
$X=-\frac{r_{o}\,p^{\hat{\phi}}}{p^{\hat{0}}},\qquad Y=\frac{r_{o}\,p^{\hat{\theta}}}{p^{\hat{0}}},\qquad b\equiv\sqrt{X^{2}+Y^{2}},$ (II.13)
which defines the impact parameter $b$ measured on the screen. Eq. (II.13) is equivalent to the common celestial-sphere parametrization by small angles $(\theta_{X},\theta_{Y})=(X/r_{o},Y/r_{o})$ at large $r_{o}$. We adopt the convention that +X+X points toward decreasing $\phi$ (hence the minus sign in $X$), and +Y+Y toward increasing $\theta$.
Figure 1: Geometry and observer-screen schematic. The Schwarzschild background with horizon at $r=2M$ (filled disk) and photon sphere at $r_{\rm ph}=3M$ (dashed circle) is shown together with a near-critical null trajectory with closest approach $r_{0}(b)$. A static observer at $r_{o}\gg M$ carries the tetrad of Eq. (II.9) and measures screen coordinates $(X,Y)$ via Eq. (II.13). The critical impact parameter $b_{c}=3\sqrt{3}\,M$ is indicated. This setup underlies the background bending integral in Eq. (III.1) and the time-dependent Born map (Eqs. (III.3)-(III.7)).
@@
Because the perturbation is axisymmetric, we may, without loss of generality, take $\theta_{o}=\pi/2$ for most calculations. At zeroth order in $\varepsilon$ the conserved energy $E=-p_{t}$ and angular momentum $L=p_{\phi}$ of the photon in the Schwarzschild background Eq. (II.1) satisfy
$\frac{X}{r_{o}}\simeq-\frac{L}{E\,r_{o}\sin\theta_{o}},\quad\frac{Y}{r_{o}}\simeq\frac{p_{\theta}}{E\,r_{o}},\quad b=\frac{L}{E}+\mathcal{O}\,\left(\frac{M}{r_{o}}\right),$ (II.14)
so that $b$ reduces to the usual ratio $L/E$ in the asymptotically flat region. In later sections, we shall describe deflection through a two-dimensional vector $\bm{\alpha}(b,t_{o})$ defined on the screen; its weak- and strong-field limits follow from the line-of-sight Born representation derived in Sec. III.
@@
$h_{\mu\nu}\rightarrow h_{\mu\nu}-\nabla_{\mu}\xi_{\nu}-\nabla_{\nu}\xi_{\mu},$ (II.15)
and on tensorial quantities derived from $h_{\mu\nu}$ accordingly. We implement two safeguards to ensure that screen observables built from Eq. (II.13) are gauge-independent at $\mathcal{O}(\varepsilon)$. First, all dynamical content of the even-parity perturbation is encoded in the Zerilli gauge-invariant master field $\Psi_{Z}$ of Eqs. (II.6)-(II.8), and the reconstruction used in Appendix A expresses $p^{\hat{a}}$ corrections in terms of $\Psi_{Z}$ and its derivatives only. Second, the tetrad is fixed operationally by the normalization and orthogonality conditions in Eq. (II.10) at the physical observer event $x_{o}^{\mu}$. With these choices, any change induced by Eq. (II.15) in the local frame is a pure $\mathcal{O}(\varepsilon)$ Lorentz transformation that leaves $(X,Y)$ invariant at that order. Concretely, one finds that the gauge variation of the screen map is a boundary term proportional to $\xi^{\mu}$ evaluated at $r_{o}$, and our asymptotic conditions impose $\xi^{\mu}=\mathcal{O}(1/r_{o})$, hence
$\delta_{\xi}X=\delta_{\xi}Y=0+\mathcal{O}\,\left(\frac{\varepsilon}{r_{o}}\right).$ (II.16)
Eq. (II.16) guarantees that the centroid motion, multi-image separations, and photon-ring modulations defined later are bona fide observables at first order.
Finally, we denote by $t_{o}$ the Schwarzschild time coordinate of the detection event $x_{o}^{\mu}$, and by $\lambda_{o}$ the corresponding affine parameter value on the photon worldline. The association $t\mapsto t_{o}$ defines the phase with which the QNM signal in Eqs. (II.6)-(II.8) is sampled by the observer and will be the sole time variable entering the deflection $\bm{\alpha}(b,t_{o})$ in Secs. III-V.
@@ III Unified Born Integral and Weak-Field Expansion
We now derive a single, time-dependent line-of-sight expression for the lensing deflection that is valid to first order in the perturbation amplitude $\varepsilon$ of Eq. (II.3) while keeping the Schwarzschild background of Eq. (II.1) exact. The resulting master Born integral will be evaluated along the unperturbed null geodesic with impact parameter $b$ and will furnish, by asymptotic expansion, both the far-field Born law (Sec. III.2) and the time-dependent strong-deflection form (Sec. IV). All observable angles are defined on the screen introduced in Sec. II.2 and are therefore gauge-safe to $\mathcal{O}(\varepsilon)$.
@@ III.1 Master time-dependent Born integral
Let $x_{0}^{\mu}(\lambda;b)$ be the null geodesic of the Schwarzschild background (Eq. (II.1)) with conserved energy $E=-p_{0t}$ and angular momentum $L=p_{0\phi}$, and let $p_{0}^{\mu}(\lambda)=dx_{0}^{\mu}/d\lambda$ be its four-momentum. The impact parameter measured at infinity is $b=L/E$ to leading order, consistent with Eq. (II.14). The background bending angle $\alpha_{0}(b)$ is given exactly by the standard integral [59]
$\alpha_{0}(b)=2\int_{r_{0}(b)}^{\infty}\frac{dr}{r^{2}}\,\left[\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}\right]^{-1/2}-\pi,$ (III.1)
@@
We switch on the first-order, axisymmetric, even-parity QNM perturbation of Eq. (II.3), encoded by the Zerilli master field $\Psi_{Z}$ that satisfies Eqs. (II.6)-(II.8). In the Born or single-scattering approximation on the curved background, the photon follows $x_{0}^{\mu}(\lambda;b)$ while its momentum acquires an $\mathcal{O}(\varepsilon)$ correction from the perturbed Hamiltonian $H=(1/2)g^{\mu\nu}p_{\mu}p_{\nu}=0$. Linearizing Hamilton’s equation $dp_{\mu}/d\lambda=-\partial_{\mu}H$ about the background solution and evaluating the source on $x_{0}^{\mu}$ and $p_{0}^{\mu}$ yields
$\delta p_{\mu}(\lambda_{o};b)=-\frac{1}{2}\int_{\lambda_{s}}^{\lambda_{o}}\,d\lambda\,\partial_{\mu}h_{\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}(\lambda)\,p_{0}^{\beta}(\lambda),$ (III.2)
where $\lambda_{o}$ and $\lambda_{s}$ denote, respectively, the values at the observer and the source, and indices are raised/lowered with the background metric. Eq. (III.2) is the covariant Born response for the photon momentum on the fixed Schwarzschild trajectory.
The approximation in Eq. (III.2) treats the perturbation to first order in $\varepsilon$ while evaluating it along the unperturbed Schwarzschild null ray. It is controlled by requiring that the induced change in the momentum remain perturbative along the integration domain, i.e., that $|\delta p_{\mu}|/|p_{0\mu}|\ll 1$ for the rays of interest. In the near-critical regime $b\to b_{c}^{+}$, the prolonged dwell time close to the photon sphere enhances the accumulated response and therefore tightens the effective bound on $\varepsilon$; this refinement is made explicit in Sec. IV when we take the near-photon-sphere limit of the same kernel.
Projecting $\delta p_{\mu}$ on the observer tetrad of Eq. (II.9) and converting to screen angles via Eq. (II.13) gives the master deflection map
$\bm{\alpha}(b,t_{o})=\bm{\alpha}_{0}(b)+\varepsilon\,\delta\bm{\alpha}(b,t_{o}),$ (III.3)
with $\bm{\alpha}_{0}(b)$ determined by Eq. (III.1), and
$\delta\alpha_{A}(b,t_{o})=\int_{\lambda_{s}}^{\lambda_{o}}\,d\lambda\,\mathcal{K}_{A}\,\left[x_{0}(\lambda);p_{0}(\lambda)\right],\quad A\in{X,Y}.$ (III.4)
@@
$\times\partial_{\mu}h_{\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}(\lambda)\,p_{0}^{\beta}(\lambda),$ (III.5)
where $e^{\nu}{}_{\hat{0}}$ is the time leg of the observer tetrad Eq. (II.9), $\Pi_{A}{}^{\ \mu}$ projects onto the $(X,Y)$ screen directions at the observer Eq. (II.13), and hats denote tetrad components at $x_{o}^{\mu}$. The above projector form is one convenient representative; Appendix B shows its equivalence to a manifestly covariant $\delta\Gamma$ line integral with cancelling boundary terms under our asymptotic conditions. Different but equivalent kernel representatives follow from integrating Eq. (III.2) by parts and using the background geodesic equation; all yield the same $\delta\bm{\alpha}$ at $\mathcal{O}(\varepsilon)$. Note that the screen deflection arises from varying $X$ and $Y$ in Eq. (II.13); keeping terms linear in $\varepsilon$ yields the projector relation $\delta\alpha_{A}=(p_{0}^{\hat{0}})^{-1}\Pi_{A}{}^{\ \hat{b}}\,\delta p_{\hat{b}}$, derived explicitly in Appendix B.
The QNM time dependence in Eqs. (II.6)-(II.8) implies that, along the background ray, the perturbation takes the form
@@
$\omega$ $=\omega_{R}+i\omega_{I}\quad\text{for}\quad\omega_{I}<0,$ (III.6)
where $t_{0}(\lambda)$ is the Schwarzschild time along $x_{0}^{\mu}(\lambda)$. Writing $t_{0}(\lambda)=t_{o}-\tau(\lambda)$ with $\tau(\lambda)\geq 0$ the light-travel delay between the encounter point and the observer event, Eq. (III.4) can be cast as
@@
$\mathcal{A}_{A}(b,\omega)=\int_{\lambda_{s}}^{\lambda_{o}}\,d\lambda\,\tilde{\mathcal{K}}_{A}\,\left[r(\lambda),\theta(\lambda)\right]\,e^{i\omega\,\tau(\lambda)},$ (III.7)
where $\tilde{\mathcal{K}}_{A}$ is the complex kernel obtained from $\mathcal{K}_{A}$ after factoring out the harmonic time dependence (the explicit form follows from Appendix B together with the RWZ reconstruction in Appendix A). Eq. (III.7) shows that the measured deflection inherits the single QNM frequency $\omega$; its amplitude $\mathcal{A}_{A}$ depends on $b$ only through the unperturbed trajectory and the delay profile $\tau(\lambda)$.
In Eq. (III.7), we interpret the QNM factor $e^{-i\omega t}$ as the ringdown component of a causal perturbation excited at finite retarded time, rather than as a globally defined normal mode on all of spacetime. Accordingly, the line-of-sight integral is to be understood under the same asymptotic assumptions used to discard boundary terms (Appendix B): effectively, the integrand has compact support or sufficiently rapid falloff along the ray, so that no additional endpoint contributions arise beyond $\mathcal{O}(\varepsilon/r_{o})$.
Eqs. (III.3)-(III.7) constitute the master, time-dependent Born formalism. Two regimes will be obtained from the same expression. In Sec. III.2 we expand Eq. (III.7) for $M/b\ll 1$ to recover the far-field Born law and the centroid wobble. In Sec. IV we analyze the contribution from segments of the trajectory that probe $r\approx 3M$, showing that Eq. (III.7) reproduces the logarithmic strong-deflection behavior once $b\to b_{c}^{+}=3\sqrt{3}M$, with the time dependence inherited from the same phase factor.
@@ III.2 Far-field (Born) expansion and centroid wobble
We now evaluate Eqs. (III.3)-(III.7) in the weak-deflection domain $M/b\ll 1$, keeping the Schwarzschild background exact along the ray but expanding the integrals in powers of $M/b$. The background bending reduces to [59, 60, 39]
$\alpha_{0}(b)=\frac{4M}{b}+\mathcal{O}\,\left(\frac{M^{2}}{b^{2}}\right),$ (III.8)
as obtained from Eq. (III.1). For the perturbation, we use Eq. (III.7) with the unperturbed trajectory approximated by a straight line of impact parameter $b$ through the asymptotically flat region, so that
$r(\lambda)\simeq\sqrt{b^{2}+z^{2}},\qquad t_{0}(\lambda)=t_{o}-\tau(\lambda),\qquad\tau(\lambda)\simeq z,$ (III.9)
where $z$ is a signed longitudinal coordinate along the asymptotic straight ray, chosen so that $z=0$ at the point of closest approach; the source and observer correspond to $z\to$-$\infty$ and $z\to+\infty$, respectively. In this approximation the delay $\tau(\lambda)$ differs from $z$ by an additive constant fixed by $\tau(\lambda_{o})=0$ and by a branch-dependent sign; these conventions only affect the overall complex normalization and can be absorbed into $c_{A}(\omega)$ in Eq. (III.10).
In the far zone the even-parity $\ell=2$ QNM behaves as an outgoing quadrupolar wave with $h_{\mu\nu}\sim r^{-1}e^{-i\omega(t-r_{*})}$, hence $\partial_{\mu}h_{\alpha\beta}\sim r^{-2}\,e^{-i\omega(t-r_{*})}$. Substituting this falloff into Eq. (III.4) or, equivalently, Eq. (III.7), we obtain the leading Born correction
$\delta\alpha_{A}(b,t_{o})=\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{c_{A}(\omega)}{b}\,\mathcal{F}_{A}(\omega b)\right\}+\mathcal{O}\,\left(\frac{M}{b^{2}}\right),\qquad A\in{X,Y},$ (III.10)
@@
$\mathcal{F}_{A}(\omega b)\equiv\int_{-\infty}^{+\infty}\,\frac{dz}{b}\,\widehat{\mathcal{K}}_{A}\,\left(\frac{\sqrt{b^{2}+z^{2}}}{b}\right)\,e^{i\omega z}$ (III.11)
is a dimensionless Fresnel-type factor encoding the finite-frequency sampling of the QNM along the nearly straight path. The explicit form of $\widehat{\mathcal{K}}_{A}$ is given in Appendix B. Two limits of Eq. (III.10) are especially transparent:
$\delta\alpha_{A}(b,t_{o})\sim\begin{cases}\displaystyle\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{\tilde{c}_{A}(\omega)}{b}\right\},&|\omega|\,b\ll 1,\\ \displaystyle\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{\hat{c}_{A}(\omega)}{b}\,e^{-|\mathrm{Im}\omega|\,b}\right\},&|\omega|\,b\gg 1,\end{cases}$ (III.12)
with $\tilde{c}_{A},\hat{c}_{A}$ calculable constants. Thus the leading time-dependent correction decays as $1/b$ and inherits the QNM phase $e^{-i\omega t_{o}}$, in agreement with the structure of Eq. (III.7).
Collecting Eqs. (III.8) and (III.10), the far-field deflection vector on the screen reads